OpenStudy (anonymous):

please help me....how do you solve the equation (x + 4)(x – 7) = -18, Eric stated that the solution would be x+4=-18=>x=-22 or (x-7)=-18=>x-11However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning

OpenStudy (anonymous):

You have two answers for x, one of them is -2 and the other is.....5

OpenStudy (anonymous):

Eric should have simplified and set the equation equal to zero first.

OpenStudy (anonymous):

i dont understand them with the => sign in it

OpenStudy (anonymous):

I think the '=>' are just supposed to be arrows.

OpenStudy (anonymous):

ok, What doesn't make sense?

OpenStudy (anonymous):

not sure what the sign means

OpenStudy (anonymous):

Did you try checking those solutions {-2, 5} in the original equation?

OpenStudy (anonymous):

not yet i will try now. It will take me a few i have a fractured right hand and it is casted and cant right well. much easier to type the write.

OpenStudy (anonymous):

"x+4=-18=>x=-22 or (x-7)=-18=>x-11" means \[\large x+4=-18 \rightarrow x=-22\] \[\large x-7=-18 \rightarrow x=-11\]

OpenStudy (anonymous):

oic ty for explaining

OpenStudy (anonymous):

You should be able to check those solutions mentally.

OpenStudy (anonymous):

I am still learning I have to write it down to be sure I am doing it right

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

when I check them they both are correct

OpenStudy (anonymous):

Bah, my mistake, they do work. Sorry. I missed one of the negative signs when I worked it out the first time. Regardless, you still need to figure out why those are the solutions.

OpenStudy (anonymous):

@zordoloom , sorry, my error, but still: could you explain how you got those answers instead of breaking the code of conduct by just giving answers?

OpenStudy (anonymous):

Do this, distribute, and then use the quadratic formula.

OpenStudy (anonymous):

(Can also factor, if you don't want to use the QF)

OpenStudy (anonymous):

true.

OpenStudy (anonymous):

Just ditribute:---> x^2-3x-10=0, then factor and you get (x+2)(x-5)=0

OpenStudy (anonymous):

This makes it easier than using the quad. form.

OpenStudy (anonymous):

k ty

OpenStudy (anonymous):

so is that wh

OpenStudy (anonymous):

Do you know how to factor?

OpenStudy (anonymous):

somewhat not real well

OpenStudy (anonymous):

That's fine. Here, distribute first. (x+4)(x-7)=18 you get (x*x+x*-7+4*x+4*-7)=-18

OpenStudy (anonymous):

When you simplify like therms you get (x^2-3x-28)=-18

OpenStudy (anonymous):

now move the -18 to the left side to make the equation equal to 0

OpenStudy (anonymous):

you now have x^2-3x-10=0

OpenStudy (anonymous):

This now factors into (x+2)(x-5)=0

OpenStudy (anonymous):

Now you have (x+2)=0 and (x-5)=0, all what you have to do is solve for x for each equatoin.

OpenStudy (anonymous):

x-5=0, move the x to the right side, x=5, do the same with the other equation, x+2=0 move the 2 to the other side, x=-2.

OpenStudy (anonymous):

You are doing simple algebra here.

OpenStudy (anonymous):

ok i see how you are doing that

OpenStudy (anonymous):

You can't just solve x+4=-18 like you did in your question. The equation must have been set to 0 first.

OpenStudy (anonymous):

ok so which equation is the right one

OpenStudy (anonymous):

I'm not sure what your question is. Since this is a second degree polynomial (an equation that has x squared when you distribute) it must have two answers. Those two answer were x=5, and x=-2